OpenStudy (anonymous):

*MEDAL 4 HELP* Which ordered pair is a solution of the equation y = -9x +4? (10 , -86) (-4, -58) (6, -41) (-6,57)

OpenStudy (anonymous):

y = -9x +4? -86=-9(10)+4 -86=-90+4 -86=-86 A